In a series resonant LCR circuit, the voltage across R is 100 volts and R = 1 k? with C = 2µF. The resonant frequency ? is 200 rad/s. At resonance the voltage across L is ?

Grade:12

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
From the given data
i=100/1000=0.1 A
VL=VC=0.1/(200*2*10-6)=250v
Thanks and Regards
Shaik Aasif
askIITians faculty
ankit singh
3 years ago
In a series resonant LCR circuit the voltage across
In a series resonant LCR circuit the voltage across R is 100 volts and R = 1k(Ω)withC=2(μ)F. The resonant frequency (ω) is 200rad/s. At resonance the voltage across L is. IXL=0.1×2500=250V.